A 73.8 kg man steps out a window and falls (from rest) 1.13 m to a sidewalk. What is his speed just before his feet strike the pavement?

If the man falls with his knees and ankles locked, the only cushion for his fall is a 0.518 cm give in the pads of his feet. Calculate the average force exerted on him by the ground in this situation. This average force is sufficient to cause cartilage damage in the joints or to break bones.
Answer #1

A 73.8 kg man steps out a window and falls (from rest) 1.13 m to a sidewalk. What is his speed just before his feet strike the pavement?

If the man falls with his knees and ankles locked, the only cushion for his fall is a 0.518 cm give in the pads of his feet. Calculate the average force exerted on ...

Here we need to apply law of conservation of energy
So at the rest position (at 1.13 m height) Potential energy is
= m*g*h
Here Mass m=73.8 Kg
Acceleration due to gravity g=9.8 m/s^2
Height h=1.13m
So potential energy=m*g*h=73.8*9.8*1.13=817.26 J


This energy will be converted to kinetic energy at the bottom most position
So m*g*h=1/2*m*v^2
So 1/2 m*v^2=817.26
v^2= (2*817.26)/73.8=22.148
V=4.7m/s

So we know that W=F*d
W= work done=817.26
F= Average force
d= distance=0.518 cm=0.00548m
So F=817.26/0.00548

F=149135.03 N